The product of the slopes of and the line perpendicular to through C is -1 in all cases. So, I can conclude that any two perpendicular lines have slopes that are negative reciprocals of each other.
1
Expert's answer
2020-10-27T18:43:43-0400

As per the given question,



The gradient of line A "=\tan \theta"

The gradient of line B"=\tan \phi"

"\theta = 90+\phi"

"\tan\theta=\tan(90+\phi)"


"=\frac{\sin(90+\phi)}{\cos(90+\phi)}"


"=\frac{\cos(\phi)}{-\sin\phi}"

"=-\cot(\phi)"

Hence, "\tan \theta. \tan \phi =-1"

Hence we can say that the product of the slope of the perpendicular lines is -1.